# VMamba: Visual State Space Model

### v3

@@ 2. Step (b)
    ($+0.0\%$, $-3$ img/s) by adjusting the CUDA implementation of the selective scan to accommodate float16 input and float32 output. This remarkably enhances the training efficiency (throughput from $165$ to $184$), despite slight speed fluctuation at test time.
@@ 5. Step (e)
    ($+0.6\%$, $+366$ img/s) by discarding the entire multiplicative branch, adopting the VSS block illustrate in Figure 3 (d), and reducing the parameter ssm-ratio (the feature expansion factor) from $2.0$ to $1.0$. This allows us to raise the layer numbers to [2,2,5,2] while reducing the FLOPs.
@@ 6. Step (f)
    ($+0.3\%$, $+161$ img/s) by reducing the parameter d_state (the SSM state dimension) from $16.0$ to $1.0$. This allows us to raise ssm-ratio back to $2.0$ and introduce the DWConv layers without increasing the FLOPs.
@@ 5 Experiments
In this section, we conduct a series of experiments to evaluate the performance of VMamba and compare it with popular benchmark models across various visual tasks. We also validate the effectiveness of the proposed 2D feature map traversing method by comparing it to alternative approaches. Additionally, we analyze the characteristics of VMamba by visualizing its effective receptive field (ERF) and activation map, and examining its scalability with increasingly long input sequences. We primarily followed the hyper-parameter settings and experimental configurations of Swin [37]. Please refer to Appendix E and F for detailed experiment settings, and Appendix H for more ablations. All experiments are conducted on a server with 8 $\times$ NVIDIA Tesla-A100 GPUs.
@@ 5.1 Image Classification
We evaluate VMamba’s performance in image classification on ImageNet-1K [9], and the comparison results against benchmark methods are summarized in Table 1. With similar FLOPs, VMamba-T achieves a top-1 accuracy of $82.6\%$, outperforming DeiT-S by $2.8\%$ and Swin-T by $1.3\%$. Notably, VMamba maintains its performance advantage at both Small and Base scale. For instance, VMamba-B achieves a top-1 accuracy of $83.9\%$, surpassing DeiT-B by $2.1\%$ and Swin-B by $0.4\%$.
In terms of computational efficiency, VMamba-T achieves a throughput of 1,686 images/s, which is either better or comparable to state-of-the-art methods. This advantage persists across VMamba-S and VMamba-B, with a throughput of $877$ images/s and $646$ images/s, respectively. Compared to SSM-based models, the throughput of VMamba-T is 1.47×1.47$\times$1.47 × higher than S4ND-Conv-T [45] and 1.08×1.08$\times$1.08 × higher than Vim-S [71], while maintaining a performance lead over them by a clear margin of $0.4\%$ and $2.1\%$, respectively.
@@ 5.2 Downstream Tasks
In this sub-section, we assess the performance of VMamba on downstream tasks, including object detection and instance segmentation on MSCOCO2017 [34], and semantic segmentation on ADE20K [70]. The training framework is developed based on the MMDetection [3] and MMSegmenation [4] library, and we follow [36] to use the Mask-RCNN [27] and UperNet [65] as the detection and segmentation network respectively.
Table 2: Left: Results of object detection and instance segmentation on MSCOCO. $AP^{b}$ and $AP^{m}$ denote box AP and mask AP, respectively. FLOPs are calculated with input size $1280\times 800$. The notation ‘1×1$\times$1 ×’ refers to models fine-tuned for $12$ epochs, and ‘3×3$\times$3 ×MS’ denotes multi-scale training for $36$ epochs. Right: Results of semantic segmentation on ADE20K. FLOPs are calculated with an input size of $512\times 2048$. ‘SS’ and ‘MS’ denote single-scale and multi-scale testing, respectively.
@@
Consistent with previous experiments, VMamba demonstrates superior performance in semantic segmentation on ADE20K with a comparable amount of parameters. As shown in Table 2, VMamba-T is $3.4\%$ mIoU higher than Swin-T and $1.9\%$ higher than ConvNeXt-T in the Single-Scale (SS) setting, and the advantage persists with Multi-Scale (MS) input. For models at the Small and Base level, VMamba-S/B outperforms NAT-S/B [26] by $2.6\%$/$2.5\%$ mIoU under the SS setting, and $1.7\%$/$1.9\%$ mIoU under the MS setting.
Figure 4: Illustration of the adaptability of VMamba to (a) downstream tasks, and (b) input image with progressively increased resolutions. Swin-T denotes Swin-T tested with scaled window sizes.
@@ Discussion
Experiment results in this subsection demonstrate VMamba’s adaptability to object detection, instance segmentation, and semantic segmentation. In Figure 4 (a), we compare VMamba’s performance with Swin and ConvNeXt, highlighting its advantages in addressing downstream tasks with comparable classification accuracy on ImageNet-1K. VMamba also exhibits greater tolerance to changes in input resolution, with linear growth in FLOPs (see Figure 4 (b)) and modest accuracy drop (further discussed in Appendix G), making it more effective and efficient compared to ViT-based methods when adapting to downstream tasks with inputs of larger spatial resolutions. This aligns with Mamba’s advanced capability in efficient long sequence modeling [17].
5.3 Analysis and Discussion
Figure 5: Illustration of the activation map for query patches indicated by red stars. Visualization results in (b) and (c) are obtained by combining the activation map of each scanning path in SS2D.
@@ 5.3 Analysis and Discussion
where $\mathbf{h_{a}}\in\mathbb{R}^{D_{k}}$ is the hidden state at step $a$, $\mathbf{M}$ denotes the temporal mask matrix of size $T\times T$ with the lower triangular part set to 1 and elsewhere 0, and $\odot$ denotes element-wise product. Detailed derivations are deferred to Appendix C.
In Eq. 4, the matrix multiplication process involving $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ closely resembles the mechanism of self-attention, despite the inclusion of $\mathbf{w}$. Concretely, the formulation of each element in $\mathbf{w}\coloneqq\left[\mathbf{w}_{1};\ldots;\mathbf{w}_{T}\right]\in\mathbb{R}^{T\times D_{k}\times D_{v}}$, i.e., $\mathbf{w}_{i}\in\mathbb{R}^{D_{k}\times D_{v}}$, can be written as $\mathbf{w}_{i}=\prod_{j=1}^{i}e^{\mathbf{A}\boldsymbol{\Delta}_{a-1+j}^{\top}}$, representing the cumulative attention weight at step $i$ computed along the scanning path.
@@
To gain an intuitive and in-depth understanding of SS2D, we further visualize the attention values in $\mathbf{Q}\mathbf{K}^{\top}$ and $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$ corresponding to a specific query patch within foreground objects (referred to as the ‘activation map’). As shown in Figure 5 (b), the activation map of $\mathbf{Q}\mathbf{K}^{\top}$ demonstrates the effectiveness of SS2D in capturing and retaining the traversed information, with all previously scanned tokens in the foreground region being activated. Furthermore, the inclusion of $\mathbf{w}$ leads to activation maps that are more focused on the neighborhood of query patches (Figure 5 (c)), which is consistent with the temporal weighting effect inherent to the formulation of $\mathbf{w}$. Nevertheless, the selective scan mechanism enables VMamba to accumulate the history along the scanning path, thereby facilitating the establishment of long-term dependencies across image patches. This is evident in the sub-figure encircled with a red box (Figure 5 (d)), where patches of the sheep far to the left (scanned in early steps) remain activated. For more visualizations and further discussion, please refer to Appendix D.
@@
The Effective Receptive Field (ERF) [41, 11] refers to the region in the input space that contributes to the activation of a certain output unit. We conduct a comparative analysis of the central pixel’s ERF across various visual backbones, both before and after training. The results presented in Figure 6 illustrate that among the models examined, only DeiT, HiViT, and VMamba, demonstrate global ERFs, whereas the others exhibit local ERFs, despite their theoretical potential for global coverage. Moreover, the linear time complexity of VMamba makes it more computationally efficient compared to DeiT and HiViT, which have quadratic costs w.r.t. the number of input patches.
Figure 6: Comparison of Effective Receptive Fields (ERF) [41] between VMamba and other benchmark models. Pixels with higher intensity indicate larger responses regarding the central pixel.
Figure 7: Performance comparison of different scanning patterns.
@@
We compare the proposed scanning pattern (i.e. Cross-Scan) to three benchmark patterns, including unidirectional scanning (Unidi-Scan), bidirectional scanning (Bidi-Scan), and cascade scanning (Cascade-Scan, scanning the data row-wise and column-wise successively). Feature dimensions are adjusted to maintain similar architectural parameters and FLOPs for fair comparison. As illustrated in Figure 7, Cross-Scan outperforms other scanning patterns in both computational efficiency and classification accuracy, highlighting its effectiveness in achieving 2D-Selective-Scan. Removing the DWConv layer, which has been observed to potentially aid the model in learning 2D spatial information, further enhances this advantage. This emphasizes the inherent strength of Cross-Scan in capturing 2D contextual information through its adoption of four-way scanning.
@@ 6 Conclusion
This paper presents VMamba, an efficient vision backbone network built with State Space Models (SSMs). VMamba integrates the benefits of selective SSMs from NLP tasks into visual data processing, bridging the gap between ordered 1D scanning and non-sequential 2D traversal through the novel SS2D module. Moreover, we have substantially improved the inference speed of VMamba through a series of architectural and implementation refinements. The effectiveness of the VMamba family has been demonstrated through extensive experiments, and the linear time complexity of VMamba makes it advantageous for downstream tasks with large-resolution inputs.
@@
While VMamba demonstrates promising experimental results, there is still room for this study to be further improved. Previous research has validated the efficacy of unsupervised pre-training on large-scale datasets (e.g., ImageNet-21K). However, the compatibility of existing pre-training methods with SSM-based architectures like VMamba, and the identification of pre-training techniques specifically tailored to such models, remain unexplored. Investigating these aspects could serve as a promising avenue for future research in architectural design. Moreover, limited computational resources have prevented us from exploring VMamba’s architecture at the Large scale, as well as conducting a fine-grained search of hyperparameters to further improve experimental performance.
@@ Appendix A Discretization of State Space Models (SSMs)
In this section, we explore the correlation between the discretized formulation of State Space Models (SSMs) obtained in Sec. 3 and those derived from the zero-order hold (ZOH) method [17], which is frequently used in studies related to SSM.
@@
$\mathbf{h}_{b}=e^{\mathbf{A}\left(\Delta_{a}+...+\Delta_{b-1}\right)}\left(\mathbf{h}_{a}+\sum^{b-1}_{i=a}\mathbf{B}_{i}u_{i}e^{-\mathbf{A}\left(\Delta_{a}+...+\Delta_{i}\right)}\Delta_{i}\right).$ (5)
@@
where $\mathbf{\overline{A_{a}}}\coloneq e^{\mathbf{A}\Delta_{a}}$ is exactly the discretized form of the evolution matrix $\mathbf{A}$ obtained by ZOH, while $\mathbf{\overline{B_{a}}}\coloneq\mathbf{B}_{a}\Delta_{a}$ is basically the first-order Taylor expansion of the discretized $\mathbf{B}$ acquired through ZOH.
@@ Appendix B Derivation of the Recurrence Relation of Selective SSMs
In this section, we derive the recurrence relation of the hidden state in selective SSMs. Given the expression of $h_{b}$ shown in Eq. 5, let us denote $e^{\mathbf{A}\left(\Delta_{a}+...+\Delta_{i-1}\right)}$ as $\mathbf{p_{A,a}^{i}}$, then its recurrence relation can be directly written as
@@
$\mathbf{p_{B,a}^{b}}$ $=e^{\mathbf{A}\left(\Delta_{a}+...+\Delta_{b-1}\right)}\sum^{b-1}_{i=a}\mathbf{B_{i}}u_{i}e^{-\mathbf{A}\left(\Delta_{a}+...+\Delta_{i}\right)}\Delta_{i}$ (8)

### v4

@@ 2. Step (b)
    ($+0.0\%$, $-3$ img/s) by adjusting the CUDA implementation of selective scan to accommodate float16 input and float32 output. This remarkably enhances the training efficiency (throughput from $165$ to $184$), despite slight speed fluctuation at test time.
@@ 5. Step (e)
    ($+0.6\%$, $+366$ img/s) by reducing the parameter ssm-ratio (the feature expansion factor) from $2.0$ to $1.0$ (also referred to as Step (d.1)), raising the layer numbers to [2,2,5,2] (also referred to as Step (d.2)), and discarding the entire multiplicative branch as illustrated in Figure 3 (c).
@@ 6. Step (f)
    ($+0.3\%$, $+161$ img/s) by introducing the DWConv layers (also referred to as Step (e.1)) and reducing the parameter d_state (the SSM state dimension) from $16.0$ to $1.0$ (also referred to as Step (e.2)), together with raising ssm-ratio back to $2.0$.
@@ 5 Experiments
In this section, we present a series of experiments to evaluate the performance of VMamba and compare it to popular benchmark models across various visual tasks. We also validate the effectiveness of the proposed 2D feature map traversal method by comparing it with alternative approaches. Additionally, we analyze the characteristics of VMamba by visualizing its effective receptive field (ERF) and activation map, and examining its scalability with longer input sequences. We primarily follow the hyperparameter settings and experimental configurations used in Swin [36]. For detailed experiment settings, please refer to Appendix E and F, and for additional ablations, see Appendix H. All experiments were conducted on a server with 8 $\times$ NVIDIA Tesla-A100 GPUs.
@@ 5.1 Image Classification
We evaluate VMamba’s performance in image classification on ImageNet-1K [9], with comparison results against benchmark methods summarized in Table 1. With similar FLOPs, VMamba-T achieves a top-1 accuracy of $82.6\%$, outperforming DeiT-S by $2.8\%$ and Swin-T by $1.3\%$. Notably, VMamba maintains its performance advantage at both Small and Base scales. For example, VMamba-B achieves a top-1 accuracy of $83.9\%$, surpassing DeiT-B by $2.1\%$ and Swin-B by $0.4\%$.
In terms of computational efficiency, VMamba-T achieves a throughput of 1,686 images/s, which is either superior or comparable to state-of-the-art methods. This advantage continues with VMamba-S and VMamba-B, achieving throughputs of $877$ images/s and $646$ images/s, respectively. Compared to SSM-based models, the throughput of VMamba-T is 1.47×1.47$\times$1.47 × higher than S4ND-Conv-T [44] and 1.08×1.08$\times$1.08 × higher than Vim-S [69], while maintaining a clear performance lead of $0.4\%$ and $2.1\%$ over these models, respectively.
@@ 5.2 Downstream Tasks
In this sub-section, we evaluate the performance of VMamba on downstream tasks, including object detection and instance segmentation on MSCOCO2017 [33], and semantic segmentation on ADE20K [68]. The training framework is based on the MMDetection [3] and MMSegmenation [4] libraries, following [35] in utilizing Mask R-CNN [26] and UperNet [63] as the detection and segmentation networks, respectively.
Table 2: Left: Results for object detection and instance segmentation on MSCOCO. $AP^{b}$ and $AP^{m}$ denote box AP and mask AP, respectively. FLOPs are calculated with an input size of $1280\times 800$. The notation ‘1×1$\times$1 ×’ indicates models fine-tuned for $12$ epochs, while ‘3×3$\times$3 ×MS’ denotes multi-scale training for $36$ epochs. Right: Results for semantic segmentation on ADE20K. FLOPs are calculated with an input size of $512\times 2048$. ‘SS’ and ‘MS’ denote single-scale and multi-scale testing, respectively.
@@
Consistent with previous experiments, VMamba demonstrates superior performance in semantic segmentation on ADE20K with a comparable amount of parameters. As shown in Table 2, VMamba-T achieves $3.4\%$ higher mIoU than Swin-T and $1.9\%$ higher than ConvNeXt-T in the Single-Scale (SS) setting, and the advantage persists with Multi-Scale (MS) input. For models at the Small and Base levels, VMamba-S/B outperforms NAT-S/B [25] by $2.6\%$/$2.5\%$ mIoU in the SS setting, and $1.7\%$/$1.9\%$ mIoU in the MS setting.
Figure 4: Illustration of VMamba’s adaptability to (a) downstream tasks and (b) input images with progressively increasing resolutions. Swin-T denotes Swin-T tested with scaled window sizes.
Figure 5: Illustration of VMamba’s resource consumption with progressively increasing resolutions. Swin-T denotes Swin-T tested with scaled window sizes.
@@ Discussion
The experimental results in this subsection demonstrate VMamba’s adaptability to object detection, instance segmentation, and semantic segmentation. In Figure 4 (a), we compare VMamba’s performance with Swin and ConvNeXt, highlighting its advantages in handling downstream tasks with comparable classification accuracy on ImageNet-1K. This result aligns with Figure 4 (b), where VMamba shows the most stable performance (i.e., modest performance drop) across different input image sizes, achieving a top-1 classification accuracy of $74.7\%$ without fine-tuning ($79.2\%$ with linear tuning) at an input resolution of $768\times 768$. While exhibiting greater tolerance to changes in input resolution, VMamba also maintains linear growth in FLOPs and memory-consumption (see Figure 5 (a) and (c)) and maintains high throughput ( Figure 5 (b)), making it more effective and efficient compared to ViT-based methods when adapting to downstream tasks with inputs of larger spatial resolutions. This aligns with Mamba’s advanced capability in efficient long sequence modeling [17].
5.3 Analysis
Figure 6: Illustration of the activation map for query patches indicated by red stars. The visualization results in (b) and (c) are obtained by combining the activation maps from each scanning path in SS2D.
@@ 5.3 Analysis
$\mathbf{y_{b}}^{(j)}=\left(\mathbf{Q_{T}}\odot\mathbf{w_{T}}^{(j)}\right){\mathbf{h_{a}}^{(j)}}+\mathbf{Q_{T}}\sum_{i=1}^{T}\left(\frac{\mathbf{w_{T}}^{(j)}}{\mathbf{w_{i}}^{(j)}}\odot\mathbf{K_{i}}\right)^{\top}\odot\left(\mathbf{V_{i}}^{(j)}\right).$ (4)
where $\mathbf{h_{a}}\in\mathbb{R}^{D_{k}}$ is the hidden state at step $a$, $\odot$ denotes element-wise product. Particularly, $\mathbf{V_{i}}^{(j)}$ is only a scalar. The formulation of each element in $\mathbf{w}\coloneqq\left[\mathbf{w}_{1};\ldots;\mathbf{w}_{T}\right]\in\mathbb{R}^{T\times D_{k}\times D_{v}}$, i.e., $\mathbf{w}_{i}\in\mathbb{R}^{D_{k}\times D_{v}}$, can be written as $\mathbf{w}_{i}=\prod_{j=1}^{i}e^{\mathbf{A}\boldsymbol{\Delta}_{a-1+j}^{\top}}$, representing the cumulative attention weight at step $i$ computed along the scanning path.
[…]
@@
where $\mathbf{M}$ denotes the temporal mask matrix of size $T\times T$ with the lower triangular part set to 1 and elsewhere 0. Please refer to Appendix C for more detailed derivations.
In Eq. 5, the matrix multiplication process involving $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ closely resembles the self-attention mechanism, despite the inclusion of $\mathbf{w}$.
@@
To gain an intuitive and in-depth understanding of SS2D, we further visualize the attention values in $\mathbf{Q}\mathbf{K}^{\top}$ and $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$ corresponding to a specific query patch within foreground objects (referred to as the ‘activation map’). As shown in Figure 6 (b), the activation map of $\mathbf{Q}\mathbf{K}^{\top}$ demonstrates the effectiveness of SS2D in capturing and retaining traversed information, with all previously scanned tokens in the foreground region being activated. Furthermore, the inclusion of $\mathbf{w}$ results in activation maps that are more focused on the neighborhood of query patches (Figure 6 (c)), which is consistent with the temporal weighting effect inherent in the formulation of $\mathbf{w}$. Nevertheless, the selective scan mechanism allows VMamba to accumulate history along the scanning path, facilitating the establishment of long-term dependencies across image patches. This is evident in the sub-figure encircled by a red box (Figure 6 (d)), where patches of the sheep far to the left (scanned in earlier steps) remain activated. For more visualizations and further discussion, please refer to Appendix D.
@@
The Effective Receptive Field (ERF) [40, 11] refers to the region in the input space that contributes to the activation of a specific output unit. We conduct a comparative analysis of the central pixel’s ERF across various visual backbones, both before and after training. The results presented in Figure 7 illustrate that among the models examined, only DeiT, HiViT, Vim and VMamba demonstrate global ERFs, while the others exhibit local ERFs despite their theoretical potential for global coverage. Moreover, VMamba’s linear time complexity enhances its computational efficiency compared to DeiT and HiViT, which incur quadratic costs w.r.t. the number of input patches. While both VMamba and Vim are based on the Mamba architecture, VMamba’s ERF is more uniform and 2D-aware than that of Vim, which may intuitively explain its superior performance.
Figure 7: Comparison of Effective Receptive Fields (ERF) [40] between VMamba and other benchmark models. Pixels with higher intensity indicate larger responses related to the central pixel.
Figure 8: Performance comparison of different scanning patterns. The proposed Cross-Scan achieves superior performance in speed while maintaining the same number of parameters and FLOPs.
@@
We compare the proposed scanning pattern (i.e. Cross-Scan) to three benchmark patterns: unidirectional scanning (Unidi-Scan), bidirectional scanning (Bidi-Scan), and cascade scanning (Cascade-Scan, scanning the data row-wise and column-wise successively). Feature dimensions are adjusted to maintain similar architectural parameters and FLOPs for a fair comparison. As illustrated in Figure 8, Cross-Scan outperforms the other scanning patterns in both computational efficiency and classification accuracy, highlighting its effectiveness in achieving 2D-Selective-Scan. Removing the DWConv layer, which has been shown to aid the model in learning 2D spatial information, further enhances this advantage. This underscores the inherent strength of Cross-Scan in capturing 2D contextual information through its adoption of four-way scanning.
@@ 6 Conclusion
This paper presents VMamba, an efficient vision backbone model built with State Space Models (SSMs). VMamba integrates the advantages of selective SSMs from NLP tasks into visual data processing, bridging the gap between ordered 1D scanning and non-sequential 2D traversal through the novel SS2D module. Furthermore, we have significantly improved the inference speed of VMamba through a series of architectural and implementation refinements. The effectiveness of the VMamba family has been demonstrated through extensive experiments, and its linear time complexity makes VMamba advantageous for downstream tasks with large-resolution inputs.
@@
While VMamba demonstrates promising experimental results, there is still room for improvement in this study. Previous research has validated the efficacy of unsupervised pre-training on large-scale datasets (e.g., ImageNet-21K). However, the compatibility of existing pre-training methods with SSM-based architectures like VMamba, as well as the identification of pre-training techniques specifically tailored for such models, remain unexplored. Investigating these aspects could serve as a promising avenue for future research in architectural design. Additionally, limited computational resources have prevented us from exploring VMamba’s architecture at the Large scale and conducting a fine-grained hyperparameter search to further enhance experimental performance. Although SS2D, the core component of VMamba, does not make specific assumptions about the layout or modality of the input data, allowing it to generalize across various tasks, the potential of VMamba for integration into more generalized tasks remains unexplored. Bridging the gap between SS2D and these tasks, along with proposing a more generalized scanning pattern for vision tasks, represents a promising research direction.
7 Acknowledgments
This work was supported by National Natural Science Foundation of China (NSFC) under Grant No.62225208 and 62406304, CAS Project for Young Scientists in Basic Research under Grant No.YSBR-117, China Postdoctoral Science Foundation under Grant No.2023M743442, and Postdoctoral Fellowship Program of CPSF under Grant No.GZB20240730.
@@ Appendix A Discretization of State Space Models (SSMs)
In this section, we explore the correlation between the discretized formulations of State Space Models (SSMs) obtained in Sec. 3 and those derived from the zero-order hold (ZOH) method [17], which is frequently used in studies related to SSMs.
@@
$\mathbf{h}_{b}=e^{\mathbf{A}\left(\Delta_{a}+\ldots+\Delta_{b-1}\right)}\left(\mathbf{h}_{a}+\sum^{b-1}_{i=a}\mathbf{B}_{i}u_{i}e^{-\mathbf{A}\left(\Delta_{a}+\ldots+\Delta_{i}\right)}\Delta_{i}\right).$ (6)
@@
where $\mathbf{\overline{A_{a}}}\coloneq e^{\mathbf{A}\Delta_{a}}$ is the exact discretized form of the evolution matrix $\mathbf{A}$ obtained by ZOH, and $\mathbf{\overline{B_{a}}}\coloneq\mathbf{B}_{a}\Delta_{a}$ represents the first-order Taylor expansion of the discretized $\mathbf{B}$ acquired through ZOH.
@@ Appendix B Derivation of the Recurrence Relation of Selective SSMs
In this section, we derive the recurrence relation of the hidden state in selective SSMs. Given the expression of $h_{b}$ shown in Eq. 6, let us denote $e^{\mathbf{A}\left(\Delta_{a}+\ldots+\Delta_{i-1}\right)}$ as $\mathbf{p_{A,a}^{i}}$. Then, its recurrence relation can be directly written as
@@
$\mathbf{p_{B,a}^{b}}$ $=e^{\mathbf{A}\left(\Delta_{a}+\ldots+\Delta_{b-1}\right)}\sum^{b-1}_{i=a}\mathbf{B_{i}}u_{i}e^{-\mathbf{A}\left(\Delta_{a}+\ldots+\Delta_{i}\right)}\Delta_{i}$ (9)